# MathOPEval: A Fine-grained Evaluation Benchmark for Visual Operations of MLLMs in Mathematical Reasoning

### v2

@@ Abstract
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
I Introduction
RECENT years have witnessed remarkable progress in visual mathematical reasoning by Multi-modal Large Language Models (MLLMs) [1, 2, 3, 4]. Given an image and a textual problem description, MLLMs generate step-by-step reasoning toward the solution. These reasoning steps often include visual operations such as annotating angles, drawing auxiliary lines, and identifying key elements (Fig. 1). Compared to text-only mathematical reasoning, this task demands stronger multi-modal capabilities, particularly in aligning textual and visual information, planning and conducting visual operations to reach the correct answer.
@@ I Introduction
Recently, using code as an intermediate representation for multi-modal mathematical reasoning has emerged as a promising approach, leading to substantial performance gains [5, 6]. In this paradigm, images are translated into corresponding Python or  codes that can accurately reconstruct them. Visual operations are then represented as code edits. Code serves as a precise and structured textual representation of images, aligning well with language-based reasoning and reducing the ambiguity often present in direct visual operations. However, recent evaluations of multi-modal mathematical reasoning have primarily focused on text-only outputs[5, 7, 8, 6]. It remains unclear how effectively current MLLMs can generate intermediate code that reflects accurate visual operations—an essential step toward precise, interpretable multi-modal mathematical reasoning (Fig. 1).
@@
Figure 2: Illustration of initial dataset for four visual operations across five visualization types. The dataset includes evaluation instructions, code, and images constructed for different tasks and visualization types. Please refer to Section III for details.
@@
Therefore, in this work, we take a first step toward evaluating multi-modal mathematical reasoning by focusing on MLLMs’ code-related capabilities in performing visual operations. We focus on two primary aspects: (1) Multi-modal Code Generation and (2) Multi-modal Code Editing. Multi-modal Code Generation evaluates models’ ability to transform the image input into the corresponding construction code. Multi-modal Code Editing encompasses three common types of visual operations: 1) Deletion tests models’ capability to identify and remove distracting elements, streamlining figures to highlight key information; 2) Modification evaluates models’ proficiency in updating images through adding auxiliary lines or altering visual elements; 3) Annotation assesses models’ ability to add numerical values, symbols, and other markings at appropriate locations, enhancing visual information expression through clear annotations.
@@
Move one step further, we manually curate a high-quality dataset. Considering the diversity of mathematical visual reasoning tasks, we carefully select five representative types of graphics: (1) geometric figures, (2) function graphs , (3) bar charts, (4) line graphs, and (5) pie charts. For each sample, we conduct detailed manual annotations to construct initial dataset as shown in Fig. 2, including: (1) original question image; (2) images after four visual operations; (3) codes and instructions for implementing these visual operations. To ensure comprehensive and reliable evaluation, we transform the initial dataset into two question types: 10,293 multiple-choice questions to test models’ selection ability of visual operations, and 7,552 open-ended questions to assess models’ ability to generate visual operations. For free-generation questions, we also establish a evaluation strategy based on Chain-of-Thought (CoT) to conduct fine-grained assessment across multiple dimensions, including content accuracy, position consistency, code correctness and completeness.
@@ II Task Formulation
For the MCE, we scrutinize multi-modal mathematical reasoning examples and thoroughly define three types of editing operations.
@@
In total, we define four evaluation tasks.
III Dataset Construction
@@
III-A Initial Dataset Construction
@@
  • Original state ($V_{\text{orig}}$, $C_{\text{orig}}$, $I_{\text{orig}}$): The original image, its code, and a textual description of the image.
@@
  • Geometric Figures: Starting with geometric problems from the Geometry3K dataset [9], we first obtain $V_{\text{orig}}$ and manually annotate the image descriptions ($I_{\text{orig}}$) and code $C_{\text{orig}}$. The deletion state involves removing specified line segments to simplify the visualization, while the modification state incorporates auxiliary geometric elements such as medians, perpendicular lines, angle bisectors, and parallel lines to facilitate problem-solving. For the annotation state, we enhance figures with point labels and line measurements from problem descriptions.
  • Function Plots: We develop a standardized approach leveraging common high school-level functions collected from the internet [^1] https://homework.study.com [^2] https://mathspace.co/us. We set the domain of the functions as [5,5][-5,5], and generate initial plots and instructions ($V_{\text{orig}}$, $C_{\text{orig}},I_{\text{orig}}$). For the deletion stage, we delete part of the domain. The modification stage implements various function transformations, while the annotation phase focuses on marking critical points, including x-axis and y-axis intersections and extreme points.
  • Statistical Charts: We consider three types of charts: bar charts, line graphs, and pie charts. We collect the original state ($V_{\text{orig}}$, $C_{\text{orig}}$, $I_{\text{orig}}$) from the ChartX dataset [10]. In the deletion state, we remove selected elements from the visualizations: bars from bar charts, points at specific x-coordinates from line graphs, and areas from pie charts. In the modification state, we adjust specific visual properties—such as changing bar heights in bar charts, altering y-coordinates of selected points in line graphs, and modifying sector proportions in pie charts. In the annotation state, we add numerical labels to represent key values: bar heights in bar charts, y-coordinates of points in line graphs, and sector percentages in pie charts.
@@
TABLE I: Basic statistics of MathOPEval.
III-B Dataset Format Conversion
Following the definition in Eq (1) and Eq (2), we transform the initial dataset $D_{\text{init}}$ into evaluation formats. Specifically, for free-form generation dataset ($D_{\text{gen}}$),
@@ III-B Dataset Format Conversion
  • MCE (Modification):
    $C_{\text{mod}}=MLLM(V_{\text{orig}},V_{\text{mod}},C_{\text{orig}},I_{\text{mod}})$
@@
III-C Dataset Statistics
The statistics of MathOPEval is shown in Table I. In the initial dataset $D_{\text{init}}$, each sample contains four sets of task-specific codes, images, and instructions. After format conversion, we obtain 7,552 questions in the generation dataset $D_{\text{gen}}$ and 10,293 questions in the multiple-choice dataset $D_{\text{mc}}$, resulting in a total of 17,845 questions. Fig. 3 shows the statistics on the proportion of different types of visualization input and lists their major context domains. We can observe that the dataset covers a diverse range of domains, including “Arts and Culture”, “Social Media and the Web”, demonstrating its broad applicability across different contexts.
IV Chain-of-Thought Evaluation Strategy
After constructing the evaluation dataset, we aim to develop an effective and efficient strategy for assessing the correctness of generated code outputs. While evaluation is straightforward in the multiple-choice format—by extracting answer choices and computing accuracy—it becomes more challenging in free-form generation settings. Given the high cost of human evaluation, we leverage state-of-the-art LLMs with strong language and coding capabilities. Specifically, we propose an automated Chain-of-Thought (CoT) evaluation strategy, utilizing the DeepSeek-V3 [11] model for reliable and unbiased code evaluation. Our evaluation process comprises two key steps: key element extraction and analytical scoring.
Key Element Extraction: To ensure precise and efficient evaluation, we instruct the evaluator model to first extract and compare key elements from both the generated and reference codes, with the specific elements varying by task type. This targeted comparison enables more reliable evaluation of code correctness and task-specific MLLM coding capabilities.
@@ IV Chain-of-Thought Evaluation Strategy
  • In MCE (Deletion), we analyze the removed elements by comparing the generated and reference $C_{out}$ against $C_{in}$.
  • In MCE (Modification), we examine the modified elements by comparing the generated and reference $C_{out}$ against $C_{in}$.
  • In MCE (Annotation), we compare the additional annotation code segments between $C_{in}$ and $C_{out}$.
Analytical Scoring: We design a comprehensive five-level scoring rule (1–2, 3–4, 5–6, 7–8, 9–10) that evaluates outputs across multiple dimensions: content accuracy, positional consistency, code correctness, and completeness. For each visualization type and task, we define detailed evaluation criteria corresponding to each score range, as outlined in the Appendix.
Aggregating the Two Steps: To illustrate the complete evaluation pipeline, Fig. 4 shows an example of our process for the MCE (Modification) task on geometric figures. Given the complete instruction, the evaluator model first identifies key elements in the generated and reference code relative to the original code. It then conducts a comprehensive analysis across the four core dimensions. Based on the defined scoring criteria, the model assigns an overall score on the code quality.
Human Evaluation: To assess the reliability of our automated evaluation, we randomly sampled 100 model-generated scores and conducted human evaluations. Results show that 92% of the scores were considered reasonable by human annotators, indicating strong alignment between automated and manual judgments. Full details of the human evaluation criteria are provided in the Appendix.
V Experiments
V-A Experimental Setup
We conduct comprehensive evaluations on a diverse set of models, including proprietary models GPT-4o [12], Qwen-VL-Max [13], and open-source models Qwen2.5-VL [13], Gemma3 [14], LLaVA-NeXT [15] Series, and specialized reasoning models QVQ-72B [13]. To minimize randomness in model outputs, we set the temperature parameter to 0. For multiple-choice tasks, we investigate various prompting strategies: (1) Direct instructs models to output answers directly; (2) CoT [16] elicits step-by-step reasoning paths; (3) Descriptive Chain-of-Thought (DCoT) [17] requires models to generate descriptions before answering; and (4) Visualization-of-Thought (VoT) [18] prompts models to imagine reasoning paths with the instruction “Visualize the state after each reasoning step.” Due to resource constraints, we evaluate Direct prompting for free-form generation tasks.
@@ V-A Experimental Setup
TABLE II: Main experimental results of four visual operations across five visualization types in multiple-choice format. “Geo” and “Func” are abbreviations for geometry and function. “AVG” represents the average accuracy across different models, visualization types, and visual operations. Bold numbers indicate the best results among all models and prompting methods, while underlined numbers denote the best accuracy for individual models across all prompting strategies.
@@
TABLE III: Main experimental results of four visual operations across five visualization types in generation format. Bold numbers indicate the best results among all models.
V-B Main Results
Tables II and III present main results for multiple-choice questions and free-form generation tasks. Our analysis of these results reveals several key findings:
Model Performance: A substantial performance gap exists between models and human baseline across all tasks. Visual operations that humans find intuitive pose significant challenges for models, with an average performance gap of 52.39% and peaking at 67.13% for geometric deletion tasks. Among the evaluated models, GPT-4o demonstrates superior performance in free-form generation with an average score of 6.08, while Qwen2.5-VL-72B-Instruct achieves the highest accuracy of 68.02% in multiple-choice questions. At the other end of the spectrum, Gemma3-4B-IT consistently shows the lowest performance, scoring 21.00% and 3.30 in multiple-choice and free-form generation respectively. The results demonstrate that generally, visual operation comprehension improves with model size, aligning with established scaling laws in language models [19].
Task Difficulty: The complexity varies significantly across different task types. In both settings, MCE (Modification) emerges as the most challenging with an average accuracy of merely 42.15% in multiple-choice questions and an average score of just 2.54 in free-form generation. MCE (Annotation), however, achieves the highest scores in both settings, indicating that while models excel at understanding and processing numerical annotations, they struggle considerably with tasks requiring detailed comprehension and manipulation of image modifications.
Visualization Challenges: Among different visualization types, function plots present the most significant challenges, achieving only 38.41% accuracy in multiple-choice setting and an average score of 2.83 in free-form generation. And the difficulty level of geometric problems follows accordingly. While pie charts and line graphs generally demonstrate better performance with average accuracies around 50%, line graphs unexpectedly show lower multiple-choice accuracy at 45.47% compared to other statistical charts, revealing specific difficulties in understanding and manipulating line graph modifications.
@@ V-C In-Depth Analysis
Reasoning-Enhanced Models vs. General Models: To investigate whether reasoning-enhanced models outperform general models of similar size in visual operations, we conduct a comparative analysis between QVQ-72B-Preview and Qwen2.5-VL-72B-Instruct. Fig. 5 reveals an interesting pattern: the reasoning-enhanced model demonstrates superior performance in MCG of mathematical functions and MCE (Deletion) of line charts, but shows comparable or slightly inferior performance across other tasks. This performance disparity suggests that while reasoning enhancement effectively improves the model’s reasoning capabilities, it does not necessarily translate to better performance in general visual operations. Such observation raises important questions about the specialization of model capabilities and the trade-offs between focused reasoning enhancement and general visual understanding.
Figure 5: Comparison between reasoning-enhanced and general-purpose models on multiple-choice format using Direct prompting strategy.
Figure 6: Analysis of shot count effects on Qwen-VL-Max’s multiple-choice performance across four visual operations in geometric figures, using CoT prompting strategy.
Impact of In-Context Learning: We investigate the effectiveness of in-context learning (ICL) using Qwen-VL-Max on geometric figures. As demonstrated in Fig. 6, ICL significantly enhances performance in MCE (Annotation) and MCE (Deletion) tasks compared to zero-shot learning, which can be attributed to the relatively straightforward nature of these operations. However, when it comes to more complex tasks such as MCG and MCE (Modification), which involve diverse geometric elements (e.g., triangles, rectangles) and various line types (including parallel lines, perpendicular lines, angle bisectors, and median lines), ICL may introduce unintended noise. This complexity and the wide variability of possible operations can actually lead to performance degradation.
@@
TABLE IV: Comparison of x-intercepts, y-intercepts, and extreme points in function-type MCE (Annotation) for multiple-choice.
@@
TABLE V: Comparison of forward and backward problems in geometric MCE (Annotation) for multiple-choice.
@@
TABLE VI: Comparison of angle and line annotation in geometric MCE (Annotation) for free-form generation.
Shot Count Analysis: Fig. 6 demonstrates that increasing the number of shots does not monotonically improve performance. Performance typically peaks at 2 shots and declines beyond 3 shots. This pattern suggests that careful selection of shot count is crucial for optimal performance. The performance degradation with higher shot counts may be attributed to two factors. On one hand, while initial shots enhance task understanding, excessive examples introduce confusion, especially with diverse configurations. On the other hand, models’ limited context window constrains their ability to process numerous shots effectively, leading to performance degradation.
Intersection Points vs. Extreme Points: Our analysis of MCE (Annotation) for function plots in Table VI reveals varying performance patterns for different mathematical features. Generally, models perform best at identifying extreme points, followed by x-intercepts, with y-intercepts proving most challenging. Interestingly, we observe that Qwen2.5-VL-3B-Instruct outperforms Qwen2.5-VL-72B-Instruct on y-intercept tasks, and Gemma3-4B-IT surpasses Gemma3-27B-IT on both types of intercepts. This counter-intuitive finding suggests that model size does not necessarily correlate with improved understanding of coordinate intersections.
Forward vs. Backward Questions: In MCE (Annotation) for geometric figures, we compare performance on forward and backward questions. The former identifies annotations based on geometric elements, e.g., what is the measure of angle ABC, while the latter identifies geometric elements based on annotations, e.g., which angle measures 120 degrees. As shown in Table VI, models consistently perform better on backward questions. We hypothesize this asymmetry might stem from the fact that backward questions provide explicit numerical anchors that help constrain the search space, while forward questions require more comprehensive geometric understanding and spatial reasoning.
@@
Points vs. Lines: Regarding the generation setting for MCE (Annotation) in geometric problems, we further categorize the annotations into two types: angle markings and line labels. As illustrated in Table VI, angle annotations demonstrate notably higher scores compared to line annotations. We attribute this performance difference to the varying complexity of placement requirements: angle markings only need to be positioned near the identified vertex, whereas line labels require to be placed at the midpoint of the line segment. This more stringent positioning requirement for line labels contributes to the increased difficulty in their accurate placement.
V-D Error Analysis

### v3

# MathOPEval: A Fine-grained Benchmark for Visual Operations of MLLMs in Mathematical Reasoning

@@ Abstract
1 Introduction
@@
Recent years have witnessed remarkable progress in visual mathematical reasoning by Multi-modal Large Language Models (MLLMs) (DBLP:journals/corr/abs-2412-15115; DBLP:journals/corr/abs-2410-21276; DBLP:journals/corr/abs-2403-05530). Given an image and a textual problem description, MLLMs generate step-by-step reasoning toward the solution. These reasoning steps often include visual operations such as annotating angles, drawing auxiliary lines, and identifying key elements (Figure 1). Compared to text-only mathematical reasoning, this task demands stronger multi-modal capabilities, particularly in aligning textual and visual information, planning and conducting visual operations to reach the correct answer.
@@
Recently, using code as an intermediate representation for multi-modal mathematical reasoning has emerged as a promising approach, leading to substantial performance gains (DBLP:conf/nips/HuSFROZSK24; DBLP:journals/corr/abs-2501-05452). In this paradigm, images are translated into corresponding Python or  codes that can accurately reconstruct them. Visual operations are then represented as code edits. Code serves as a precise and structured textual representation of images, aligning well with language-based reasoning and reducing the ambiguity often present in direct visual operations. However, recent evaluations of multi-modal mathematical reasoning have primarily focused on text-only outputs (DBLP:conf/nips/HuSFROZSK24; DBLP:journals/corr/abs-2412-12932; DBLP:journals/corr/abs-2501-05452). It remains unclear how effectively current MLLMs can generate intermediate code that reflects accurate visual operations—an essential step toward precise, interpretable multi-modal mathematical reasoning (Figure 1).
@@
Figure 2: Illustration of initial dataset for four visual operations across five visualization types. The dataset includes evaluation instructions, code, and images constructed for different tasks and visualization types. Please refer to Section 3 for details.
@@
Therefore, in this work, we take a first step toward evaluating multi-modal mathematical reasoning by focusing on MLLMs’ code-related capabilities in performing visual operations. We focus on two primary aspects: (1) Multi-modal Code Generation and (2) Multi-modal Code Editing. Multi-modal Code Generation evaluates models’ ability to transform the image input into the corresponding construction code. Multi-modal Code Editing encompasses three common types of visual operations: 1) Deletion tests models’ capability to identify and remove distracting elements, streamlining figures to highlight key information; 2) Modification: evaluates models’ proficiency in updating images through adding auxiliary lines or altering visual elements; 3) Annotation assesses models’ ability to add numerical values, symbols, and other markings at appropriate locations, enhancing visual information expression through clear annotations.
@@
Move one step further, we manually curate a high-quality dataset. Considering the diversity of mathematical visual reasoning tasks, we carefully select five representative types of graphics: (1) geometric figures, (2) function graphs , (3) bar charts, (4) line graphs, and (5) pie charts. For each sample, we conduct detailed manual annotations to construct initial dataset as shown in Figure 2, including: (1) original question image; (2) images after four visual operations; (3) codes and instructions for implementing these visual operations. To ensure comprehensive and reliable evaluation, we transform the initial dataset into two question types: 10,293 multiple-choice questions to test models’ selection ability of visual operations, and 7,552 open-ended questions to assess models’ ability to generate visual operations. For free-generation questions, we also establish a evaluation strategy based on Chain-of-Thought (CoT) to conduct fine-grained assessment across multiple dimensions, including content accuracy, position consistency, code correctness and completeness.
@@ 2 Task Formulation
For the MCE, we scrutinize multi-modal mathematical reasoning examples and thoroughly define three types of editing operations. In total, we define four evaluation tasks.
3 Dataset Construction
@@
3.1 Initial Dataset Construction
@@
  • Original state ($V_{\text{orig}}$, $C_{\text{orig}}$, $I_{\text{orig}}$): The original image, its code, and a textual description.
@@
  • Geometric Figures: Starting with geometric problems from the Geometry3K dataset (DBLP:conf/acl/LuGJQHLZ20), we first obtain $V_{\text{orig}}$ and manually annotate the image descriptions ($I_{\text{orig}}$) and code $C_{\text{orig}}$. The deletion state involves removing specified line segments to simplify the visualization, while the modification state incorporates auxiliary geometric elements such as medians, perpendicular lines, angle bisectors, and parallel lines to facilitate problem-solving. For the annotation state, we enhance figures with point labels and line measurements from problem descriptions.
  • Function Plots: We develop a standardized approach leveraging common high school-level functions collected from the internet [^1] https://homework.study.com [^2] https://mathspace.co/us. We set the domain of the functions as [5,5][-5,5], and generate initial plots and instructions ($V_{\text{orig}}$, $C_{\text{orig}},I_{\text{orig}}$). For the deletion stage, we delete part of the domain. The modification stage implements various function transformations and the annotation phase focuses on marking critical points, including x-axis and y-axis intersections and extreme points.
  • Statistical Charts: We consider three types of charts: bar charts, line graphs, and pie charts. We collect the original state ($V_{\text{orig}}$, $C_{\text{orig}}$, $I_{\text{orig}}$) from the ChartX dataset (DBLP:journals/corr/abs-2402-12185). In the deletion state, we remove selected elements from the visualizations: bars from bar charts, points at specific x-coordinates from line graphs, and areas from pie charts. In the modification state, we adjust specific visual properties—such as changing bar heights in bar charts, altering y-coordinates of selected points in line graphs, and modifying sector proportions in pie charts. In the annotation state, we add numerical labels to represent key values: bar heights in bar charts, y-coordinates of points in line graphs, and sector percentages in pie charts.
3.2 Dataset Format Conversion
Following the definition in Eq 1 and Eq 2, we transform the initial dataset $D_{\text{init}}$ into evaluation formats. Specifically, for free-form generation dataset ($D_{\text{gen}}$),
@@ 3.2 Dataset Format Conversion
Figure 4: Basic statistics of MathOPEval.
@@
  • MCE (Modification): $C_{\text{mod}}=MLLM(V_{\text{orig}},V_{\text{mod}},C_{\text{orig}},I_{\text{mod}})$
@@
3.3 Dataset Statistics
The statistics of MathOPEval is shown in Table 4. In the initial dataset $D_{\text{init}}$, each sample contains four sets of task-specific codes, images, and instructions. After format conversion, we obtain 7,552 questions in the generation dataset $D_{\text{gen}}$ and 10,293 questions in the multiple-choice dataset $D_{\text{mc}}$, resulting in a total of 17,845 questions. Figure 4 shows the statistics on the proportion of different types of visualization input and lists their major context domains. We can observe that the dataset covers a diverse range of domains, including “Arts and Culture”, “Social Media and the Web”, demonstrating its broad applicability across different contexts.
4 Chain-of-Thought Evaluation Strategy
After constructing the evaluation dataset, we aim to develop an effective and efficient strategy for assessing the correctness of generated code outputs. While evaluation is straightforward in the multiple-choice format—by extracting answer choices and computing accuracy—it becomes more challenging in free-form generation settings. Given the high cost of human evaluation, we leverage state-of-the-art LLMs with strong language and coding capabilities. Specifically, we propose an automated Chain-of-Thought (CoT) evaluation strategy, utilizing the DeepSeek-V3 (deepseekv3) model for reliable code evaluation. Our evaluation process comprises two key steps: key element extraction and analytical scoring.
Key Element Extraction To ensure precise and efficient evaluation, we instruct the evaluator model to first extract and compare key elements from both the generated and reference codes, with the specific elements varying by task type. This targeted comparison enables more reliable evaluation of code correctness and task-specific MLLM coding capabilities.
@@ 4 Chain-of-Thought Evaluation Strategy
  • In MCE, we analyze the differences between the generated output and the reference output by comparing both versions of $C_{out}$ against the original input $C_{in}$.
Analytical Scoring We design a comprehensive five-level scoring rule (1–2, 3–4, 5–6, 7–8, 9–10) that evaluates outputs across multiple dimensions: content accuracy, positional consistency, code correctness, and completeness. For each visualization type and task, we define detailed evaluation criteria corresponding to each score range, as outlined in the Appendix.
Human Evaluation To assess the reliability of our automated evaluation, we randomly sampled 100 model-generated scores and conducted human evaluations. Results show that 92% of the scores were considered reasonable by human annotators, indicating strong alignment between automated and manual judgments. Full details of the human evaluation criteria are provided in the Appendix.
5 Experiments
5.1 Experimental Setup
We conduct evaluations on a diverse set of models, including proprietary models GPT-4o (DBLP:journals/corr/abs-2410-21276), Qwen-VL-Max (DBLP:journals/corr/abs-2412-15115), and open-source models Qwen2.5-VL (DBLP:journals/corr/abs-2412-15115), Gemma3 (DBLP:journals/corr/abs-2403-05530), LLaVA-NeXT (DBLP:conf/cvpr/LiuLLL24) Series, and specialized reasoning model QVQ-72B (DBLP:journals/corr/abs-2412-15115). To minimize the randomness, we set the temperature parameter to 0. For multiple-choice tasks, we investigate various prompting strategies: (1) Direct instructs models to output answers directly; (2) CoT (DBLP:conf/nips/Wei0SBIXCLZ22) elicits step-by-step reasoning paths; (3) Descriptive Chain-of-Thought (DCoT) (DBLP:journals/corr/abs-2311-09193) requires models to generate descriptions before answering; and (4) Visualization-of-Thought (VoT) (DBLP:conf/nips/WuMZ [^0] W24) prompts models to imagine reasoning paths with the instruction “Visualize the state after each reasoning step.” Due to resource constraints, we evaluate Direct prompting for free-form generation tasks.
5.2 Main Results
Tables 1 and 2 present main results for multiple-choice questions and free-form generation tasks. Our analysis of these results reveals several key findings:
Model Performance: A substantial performance gap exists between models and human baseline across all tasks. Visual operations that humans find intuitive pose significant challenges for models, with an average performance gap of 52.39% and peaking at 67.13% for geometric deletion tasks. Among the evaluated models, Qwen2.5-VL-72B-Instruct achieves the highest accuracy of 68.02% in multiple-choice questions, while GPT-4o demonstrates superior performance in free-form generation with an average score of 6.08. At the other end of the spectrum, Gemma3-4B-IT consistently shows the lowest performance, scoring 21.00% and 3.30 in multiple-choice and free-form generation respectively. The results demonstrate that generally, visual operation performance improves with model size, aligning with established scaling laws in language models (DBLP:journals/corr/abs-2001-08361).
@@ 5.2 Main Results
Table 1: Main experimental results of four visual operations across five visualization types in multiple-choice format. “Geo” and “Func” are abbreviations for geometry and function. “AVG” represents the average accuracy across different models, visualization types, and visual operations. Bold numbers indicate the best results among all models and prompting methods, while underlined numbers denote the best accuracy for individual models across all prompting strategies.
@@
Table 2: Main experimental results of four visual operations across five visualization types in generation format. Bold numbers indicate the best results among all models.
Task Difficulty: The complexity varies significantly across different task types. In both settings, MCE(Modification) task emerges as the most challenging with an average accuracy of merely 42.15% in multiple-choice questions and an average score of just 2.54 in free-form generation. MCE(Annotation), however, achieves the highest scores in both settings, indicating that while models excel at understanding and processing numerical annotations, they struggle considerably with tasks requiring detailed comprehension and manipulation of image modifications.
Visualization Challenges: Among different visualization types, function plots present the most significant challenges, achieving only 38.41% accuracy in multiple-choice setting and an average score of 2.83 in free-form generation. And the difficulty level of geometric problems follows accordingly. While bar and pie charts demonstrate better performance with average accuracies around 50%, line graphs unexpectedly show lower multiple-choice accuracy at 45.47% compared to other statistical charts, revealing specific difficulties in understanding and manipulating line graph.
@@ 5.3 In-Depth Analysis
Reasoning-Enhanced Models vs. General Models: To investigate whether reasoning-enhanced models outperform general models of similar size in visual operations, we conduct a comparative analysis between QVQ-72B-Preview and Qwen2.5-VL-72B-Instruct. Figure 5 reveals an interesting pattern: the reasoning-enhanced model demonstrates superior performance in MCG of mathematical functions and MCE(Deletion) of line charts, but shows comparable or slightly inferior performance across other tasks. This performance disparity suggests that while reasoning enhancement effectively improves the model’s reasoning capabilities, it does not necessarily translate to better performance in general visual operations. Such observation raises important questions about the specialization of model capabilities and the trade-offs between focused reasoning enhancement and general visual understanding.
Figure 5: Comparison between reasoning-enhanced and general-purpose models on multiple-choice format using Direct prompt.
Figure 6: Analysis of shot count effects on Qwen-VL-Max’s multiple-choice performance across four visual operations using CoT prompt.
Impact of In-Context Learning: We investigate the effectiveness of in-context learning (ICL) using Qwen-VL-Max on geometric figures. As demonstrated in Figure 6, ICL significantly enhances performance in MCE(Annotation) and MCE(Deletion) compared to zero-shot learning, which can be attributed to the relatively straightforward nature of these operations. However, when it comes to more complex tasks such as MCG and MCE(Modification), which involve diverse geometric elements (e.g., triangles, rectangles) and various line types (e.g., parallel, perpendicular, median lines and angle bisectors), ICL may introduce unintended noise. This complexity and the wide variability of possible operations can actually lead to performance degradation.
Shot Count Analysis: Figure 6 demonstrates that increasing the number of shots does not monotonically improve performance. We observe that performance generally peaks at two shots and begins to decline with three or more. This highlights the importance of carefully tuning the shot count for optimal results. We attribute this performance degradation to two primary factors. On one hand, while initial examples help the model grasp the task, an excess of them can introduce noise or conflicting signals, especially with diverse inputs. On the other hand, models’ finite context window inherently limits its ability to process numerous examples, leading to information loss and diminished performance.
Intersection Points vs. Extreme Points: In Table 4 and 4, models generally perform best at identifying extreme points, followed by x-intercepts, with y-intercepts proving most challenging. Interestingly, we observe that Qwen2.5-3B-Instruct outperforms its 72B version on y-intercept tasks, and Gemma3-4B-IT surpasses its 27B version on all types. This counter-intuitive finding suggests that model size does not necessarily correlate with improved understanding of coordinate intersections.
Model X Y Extreme AVG
GPT-4o 6.03 6.88 4.71 5.87
Qwen-VL-Max 5.26 6.50 4.06 5.27
Qwen2.5-VL-3B-Instruct 3.88 4.44 4.20 4.17
Qwen2.5-VL-72B-Instruct 5.16 6.52 4.98 5.55
Gemma3-4B-IT 1.01 1.01 1.00 1.01
Gemma3-27B-IT 1.00 1.00 1.00 1.00
LLaVA-NeXT-8B 4.04 3.33 4.32 3.89
LLaVA-NeXT-72B 4.22 4.40 3.40 4.00
AVG 3.82 4.26 3.46 3.85
Table 3: Analysis of identifying function characteristics by MCE(Annotation) in free-form generation format.
@@
Table 4: Analysis of identifying function characteristics by MCE(Annotation) in multiple-choice format.
Forward vs. Backward Questions: As shown in Table 6, we compare performance on forward and backward questions in MCE(Annotation). The former identifies annotations based on geometric elements, e.g., what is the measure of angle ABC, while the latter identifies geometric elements based on annotations, e.g., which angle measures 120 degrees. We observe that models consistently perform better on backward questions. We hypothesize this asymmetry might stem from the fact that backward questions provide explicit numerical anchors that help constrain the search space, while forward questions require more comprehensive geometric understanding and spatial reasoning.
@@
Table 5: Comparison of forward and backward problems in geometric MCE(Annotation) for multiple-choice format.
@@
Table 6: Comparison of angle and line in geometric MCE(Annotation) for free-form generation.
@@
Points vs. Lines: Regarding the generation setting for MCE(Annotation) in geometric problems, we further categorize the annotations into two types: angle markings and line labels. As illustrated in Table 6, angle annotations demonstrate notably higher scores compared to line annotations. We attribute this performance difference to the varying complexity of placement requirements: angle markings only need to be positioned near the identified vertex, whereas line labels require to be placed at the midpoint of the line segment. This more stringent positioning requirement for line labels contributes to the increased difficulty in their accurate placement.
5.4 Error Analysis